Phosphorus pentachloride pcl5 dissociate to pcl3 and cl2 .at 200ºc and 1.600atm total pressure, equilibrium is reached when 35.5% of pcl5 are dissociated . calculate
a) the constant kp balance
b) the change in standard Free energy of reaction data R=8.31jk-1
1
Expert's answer
2019-12-03T08:25:29-0500

1.      PCl5= PCl3 +Cl

"Kp=Kc (RT)^n"

Kc of the reaction is 1, the number of moles n resulting is 2 and the temperature is (273.15+35.5)=308.65K

Therefore, Kp=1(8.31*308.65)=6.57*10^6

2.      The standard free energy"SFe=nRT"

SFe=1 *8.31*308.65= 2564.88J